OpenStudy (anonymous):

How man moles, molecules and atoms are there in 1.0L of water?

OpenStudy (lgbasallote):

if i remember this right (feel free to correct me)... \[\large 1.0 \text{L} H_2 O \times \frac{1 \text{mole} H_2 O}{1\text{L} H_2 O} \times \frac{6.022 \times 10^{23} \text{molecules} H_2 O}{1 \text{mole} H_2 O} \times \frac{ 1 \text{atom} H_2 O}{1 \text{molecule} H_2 O}\]